The perimeter of a rectangle is the expression (6x+10). Find one possible algebraic expression for the length and one possible algebraic expression for the width
Answer by oscargut(2103)

perimeter = 2 (L*W) = 6x+10 = 2(3x+5)
Answer: Width can be 5 and length 3x

I can help you with all of your questions at: mthman@gmail.com
Thanks